A manufacturing plant conducted a survey to determine it's employees reaction toward a proposed change in working hours.a breakdown of the response is as follows: production:17 agree 23 disagree. Office:8 agree 2 disagree. Suppose an employee is chosen at random with events being defined as A:the employee works in the production team. B:the employee agreed with the proposed change. C: the employee works in the office. D: the employee disagree with the proposed change. 4.2.1) P(A). 4.2.2) P(A or B). 4.2.3)P(C and D). 4.2.4)P(c/D)
1
Expert's answer
2020-04-20T11:14:50-0400

4.2.1) "P(A)=\frac{Production\ employees}{Totalemployees}=\frac{40}{50}=0.8"

"\bold{P(A)=\frac{40}{50}=0.8}"


4.2.2) "P(A\cup B )=P(A)+P(B)-P(A\cap B)"

"P(A\cup B )=P(A)+P(B)-P(A|B)P(B)"

where,

"P(B)=\frac{Agreed\ employees}{Total\ employees}=\frac{25}{50}=0.5\\
P(A|B)=\frac{Agreed\ production\ employees}{Agreed\ Employees}=\frac{17}{25}=0.68\\"


"P(A\cup B )=0.8+0.5-0.68*0.5\\
P(A\cup B )=0.96\\
\bold{P(A\ or\ B )=0.96}"


4.2.3) "P(C \cap D)=\frac{Disagreed\ office\ employees\ }{Total\ employees}\\"

"P(C \cap D)=\frac{2}{50}\\
P(C\cap D)=0.04\\
\bold{P(C\ and\ D)=0.04}"


4.2.3) "P(C|D)=\frac{P(C\cap D)}{P(D)}"

where,

"P(D)=\frac{Disgreed\ employees}{Total\ employees}=\frac{25}{50}=0.5\\"

"P(C|D)=\frac{0.04}{0.5}\\
\bold{P(C|D)=0.08}"